the atomic mass of an element is known to be 32,064u. one isotope occurs naturally with an abundance of 95.018% and has atomic mass of 31.971. If there is one other main isotope of this element, what will be its mass number?
1
Expert's answer
2019-10-25T05:30:37-0400

Atomic mass of an element is 32.064u.

One isotope has abundance of 95.018% and has atomic mass 31.971u.

The other isotope has abundance of 4.982 and let its atomic mass be x.

Now,

"(95.018\u00d731.971 + x\u00d74.982)\/100 = 32.064"

"3037.82048 + 4.982x = 3206.4"

"x=(3206.4-3037.82048)\/4.982"

"x=33.8377198(Answer)"

The mass number of other isotope is 33.8377198u